Which of the following statements are true and which are false? Please give reasons for the answers.
1.) The equation r=acos(θ+α)+bsin(θ+α) represents a circle.
2.) The direction ratios of the line x-5=5-y, z=5 are 1,1,5.
3.) The intersection of a plane and a cone can be a pair of lines.
4.) The angle between the planes x+2y+2z=5 and 2x+2y+3=0 is 60º
5.) The equations 2x²+y²+3z²+4x+4y+18z+34=0, 2x²-y²=4y-4y-4x represent a real conic.
1
Expert's answer
2019-07-21T17:35:05-0400

1) True

"cosine \u00a0 and \u00a0sine \u00a0are \u00a0periodic, so\u00a0(\\theta+\\alpha) \u00a0will\u00a0 represent\u00a0same\u00a0figure\u00a0as \u00a0\\theta,\\\\\u00a0turned\u00a0 by\u00a0 some\u00a0 angle \\\\ \u00a0\\\\\npolar\u00a0coordinates\u00a0is: x=rsin\\phi, \u00a0y=rcos\\phi \\\\ so,\u00a0in \u00a0Cartesian\u00a0coordinates\u00a0: \\\\\u00a0r=\\sqrt{x^2+y^2}, \u00a0cos\\theta = \\frac{y}{r} = \\frac{y}{\\sqrt{x^2+y^2}},\u00a0 sin\\theta=\\frac{x}{r} = \\frac{x}{\\sqrt{x^2+y^2}} \\\\\u00a0\\\\ our \u00a0equation:r=acos(\\theta+\\alpha) + bsin(\\theta+\\alpha) \\\\ same\u00a0figure: r=acos(\\theta) + bsin(\\theta) \\\\\\sqrt{x^2+y^2} = a\\frac{y}{\\sqrt{x^2+y^2}} + b\\frac{x}{\\sqrt{x^2+y^2}} \\\\\u00a0\\\\\nx^2+y^2 = ax+by \\\\x^2-ax+\\frac{a^2}{4} + y^2-by+\\frac{b^2}{4} - \\frac{a^2}{4} - \\frac{b^2}{4} = 0\\\\ \u00a0\\\\\n(x-\\frac{a}{2})^2 + (y-\\frac{b}{2})^2 = \\frac{a^2+b^2}{4} \\\\\u00a0\\\\ we\u00a0got \u00a0circle\u00a0 equation"



2)False.

x-5 = 5-y, so y = -x + 10 and in xOy plane directions are {1: -1} or\and {-1: 1}, so, {1:1:5} doesnt fit.


3)False, conic section can be parabola, hyperbola. ellipse and line


4)False

our planes: x+2y+2z=5 and 2x+2y+3=0

we have 2 normal vectors: a=(1,2,2) and b=(2,2,0)


the angle X can be found by the formula cos(X) = (a, b) \ (||a|| * ||b||)

so,

"cos(X) = \\frac{2+4}{\\sqrt{1+4+4}\\sqrt{4+4}} = \\frac{6}{3\\sqrt{8}} = \\frac{\\sqrt{2}\\sqrt{2}}{2\\sqrt{2}} = \\frac{\\sqrt{2}}{2}, so\u00a0angle \u00a0 is\u00a0 45\u00a0degrees"



5)False

"2x\u00b2+y\u00b2+3z\u00b2+4x+4y+18z+34=0 \\\\ \u00a0\\\\(2x^2+4x+2) + (2y^2+4y+2) +(3z^2+18z+27) +3 = 0 \\\\\u00a0\\\\\n(\\sqrt{2}x+\\sqrt{2})^2+(\\sqrt{2}y+\\sqrt{2})^2 + (\\sqrt{3}z+3\\sqrt{3})^2 = -3 \u00a0\u00a0\u00a0\\implies no\u00a0solutions"





"2x\u00b2-y\u00b2=4y-4y-4x \\\\\u00a0\\\\(2x^2+4x+2) - (y^2) - 2= 0 \\\\ 2 = (\\sqrt{2}x + \\sqrt{2})^2 - y^2 \\\\ \u00a0\\\\ 1 = \\frac{(\\sqrt{2}x + \\sqrt{2})^2}{2} - \\frac{y^2}{2} \\implies its\u00a0hyperbola"